OpenStudy (anonymous):

The Population of Jamesburg is 12,000. Junction city's population is 3/25 of Jamesburg's. What is Junction City's population?

OpenStudy (anonymous):

Formulka to be used is: Population of Junction city = 3/25 times the population of Jamesburg Population of Junction city = 3/25 *12,000 =3*480 =1440

OpenStudy (zaarin):

1440